# The Coral Protocol Open Infrastructure Connecting The Internet of Agents

### v1

@@ Abstract
The Coral Protocol is an open and decentralized collaboration infrastructure that enables communication, coordination, trust and payments for The Internet of Agents. It addresses the growing need for interoperability in a world where organizations are deploying multiple specialized AI agents that must work together across domains and vendors. As a foundational platform for multi-agent AI ecosystems, Coral establishes a common language and coordination framework allowing any agent to participate in complex workflows with others. Its design emphasizes broad compatibility, security, and vendor neutrality, ensuring that agent interactions are efficient and trustworthy. In particular, Coral introduces standardized messaging formats for agent communication, a modular coordination mechanism for orchestrating multi-agent tasks, and secure team formation capabilities for dynamically assembling trusted groups of agents. Together, these innovations position Coral Protocol as a cornerstone of the emerging “Internet of Agents,” unlocking new levels of automation, collective intelligence, and business value through open agent collaboration.
@@ 1   Introduction
  • Dynamic Agent Discovery and Capability Registration: Agents can advertise their capabilities and discover others through standardized mechanisms, allowing seamless composition of multi-agent workflows without hardcoded integrations or platform-specific logic.
  • Secure Team Formation and Workflow Coordination: Coral enables the on-demand assembly of agent teams with authenticated identities, assigned roles, and controlled data access. Teams can collaboratively execute complex tasks while preserving privacy, trust, and auditability via a secure coordination layer and optional blockchain-based logging.
@@ 2   Enabling Concepts
In this section, we review several key concepts that underpin the Coral Protocol’s vision of interoperable AI agents. We cover recent progress in large language models, autonomous AI agents and their use of tools, the emergence of model context protocols, and the advent of multi-agent AI systems. Together, these advances set the stage for why a unifying interoperability protocol is needed.
@@ 2.4   Model Context Protocol (MCP)
2.5   Multi–AI Agent Systems
@@
There are several motivations for using multiple AI agents together. One is specialization: one agent might be an expert in math, another in coding, another in interacting with humans. By having them communicate, each sub-problem can be handled by the best-suited agent. For example, in a software development assistant, one agent could generate code while another agent reviews it for errors – akin to a pair programming scenario. Another motivation is parallelism: agents can work on different subtasks simultaneously and then share results, speeding up complex workflows. Yet another is emergent behavior: sometimes groups of agents can exhibit intelligent behaviors that single agents cannot, by bouncing ideas off each other or negotiating. A recent study by Tran et al. (2025) notes that LLM-based multi-agent systems enable "groups of intelligent agents to coordinate and solve complex tasks collectively at scale, transitioning from isolated models to collaboration-centric approaches" [18]. In other words, we’re beginning to move from thinking about one AI in isolation to networks of AIs collaborating, which some see as steps toward an "artificial collective intelligence."
@@ 2.6   Multi AI Agent Collaboration
Several proof-of-concept systems have explored multi-agent collaboration. The CAMEL framework (2023) showed that two LLM agents can role-play as a "user" and an "assistant" to iteratively solve a task together, essentially allowing the AI to have a conversation with itself from different perspectives to refine a solution. Microsoft’s AutoGen [20] provides a programming framework for composing *conversational* interactions among multiple agents (and humans), so that one can build workflows where agents ask each other for help [20]. For instance, one agent could be tasked with decomposing a problem into steps, then delegating each step to other specialized agents, and finally aggregating the results. Experiments with AutoGen found that such setups can handle complex queries in domains like coding or decision-making more effectively than a single agent working alone [20]. Another striking example is the Generative Agents simulation mentioned earlier, where 25 agents inhabited a small virtual town – here the focus was on agents having social interactions with each other (e.g. sharing information, forming plans to organize a party) and it demonstrated that coherent multi-agent dynamics can emerge from relatively simple principles plus consistent language-model reasoning [15].
@@
This is where an interoperability framework becomes crucial: it can provide a structured medium for agent interaction (e.g., a shared memory or a common set of message types). The Coral Protocol is positioned in this landscape as an infrastructure to help organize and facilitate these multi-agent interactions. It extends ideas from MCP so that agents not only access data/tools uniformly, but also discover and communicate with each other in a standardized way. By having a common protocol, an "agent society" can form where each agent knows how to announce its capabilities, listen for requests, and share results in a mutual language. This would allow, for example, an agent built by Company A to collaborate with an agent from Company B if both speak Coral/MCP, much as devices on the internet interoperate via TCP/IP. In summary, multi-agent AI systems are an exciting and fast-evolving area — moving from single, siloed AI assistants towards networks of cooperating agents. The recent progress in this domain (just in the past couple of years) underscores the need for interoperability solutions: to manage complexity, avoid reinventing integration logic, and unlock the full potential of collective AI intelligence. The Coral Protocol’s goal of facilitating agent interoperability directly addresses this need, aiming to enable robust, framework-agnostic collaboration among AI agents in the coming "society of AI agents."
@@ 2.8   Blockchains and Secure Payments
Blockchains provide a decentralized, tamper-evident ledger for recording all financial exchanges among agents. By design, every transaction written to the chain is immutable and publicly verifiable, ensuring reliable auditability. In the Coral Protocol, agent payments (and any inter-agent transfers) are committed to the blockchain, creating an indelible audit trail of who paid whom and when. This means that all payments for services are transparently recorded and cannot be forged or altered after the fact.
@@ 3.1.1 Agent-to-Agent (A2A) Protocol
Google introduced the Agent-to-Agent (A2A) protocol, an open, vendor-neutral standard designed to enable seamless communication and collaboration among AI agents across various platforms and ecosystems [17]. A2A focuses on task-oriented interactions, providing standardized task objects with clearly defined lifecycles, and supporting the exchange of task-related artifacts to facilitate collaborative workflows. Built on established web standards like HTTP, JSON-RPC, and Server-Sent Events (SSE), the protocol ensures compatibility with existing technologies while offering enterprise-grade security through robust authentication and authorization aligned with OpenAPI. Additionally, A2A supports dynamic discovery of agent capabilities via JSON-based "agent cards" and handles both short-term tasks and long-running processes with real-time feedback. Google’s collaboration with over 50 industry partners—including Salesforce, SAP, and Workday—highlights the broad industry commitment toward achieving standardized agent interoperability.
@@
In practice, A2A is poised to facilitate complex workflows in both enterprise and open-source contexts. In enterprises, multiple AI agents could coordinate processes across business platforms that typically don’t talk to each other today. For example, an agent integrated with a CRM like Salesforce might invoke another agent on an ERP system such as SAP to automatically fulfill an order or update financial records [6]. Organizations are exploring scenarios like a sales assistant agent requesting a finance agent to generate a pricing quote, or a customer support chatbot querying an inventory management agent for stock availability – all through A2A’s standardized agent-to-agent calls without human middleware [7]. Because A2A is an open and framework-neutral protocol, it also lends itself to open-source innovation. Developers and researchers can orchestrate modular AI agents (for instance, linking a data analysis agent with a visualization agent) using A2A as the lingua franca for inter-agent dialogue. Indeed, popular toolkits like LangChain have begun integrating A2A, allowing independent or specialized agents to be composed into larger AI workflows in research and development settings.
@@
Despite its promise, the A2A protocol has some limitations that motivate the development of more comprehensive multi-agent frameworks. Notably, A2A focuses on the mechanics of messaging and doesn’t define a shared ontology or common knowledge base for semantics – agents still must understand the content of messages (often plain text or JSON) based on their own models or agreements, which can limit out-of-the-box interoperability in complex domains. The protocol is also in an early stage (initially released as a draft specification in 2025), so its extensibility to all possible use cases is still evolving and industry adoption is just beginning. Moreover, A2A by itself addresses communication rather than higher-level cognition or coordination among agents. It enables agents to talk but does not inherently provide collective planning, sophisticated negotiation, or “hive-mind” reasoning capabilities. Implementing such intelligence requires additional layers on top of A2A – for example, orchestration frameworks like Google’s Agent Development Kit (ADK) are intended to manage workflows and decision logic among multiple agents using the A2A channel. In summary, A2A lays an important foundation for agent interoperability, but it remains a low-level protocol; achieving robust multi-agent collaboration will also require tackling semantic standards, advanced coordination strategies, and other gaps that A2A alone does not fill.
@@ 3.1.3 AGNTCY (Internet of Agents Initiative)
Together, these pieces enable what AGNTCY envisions: for example, an organization could deploy multiple AI agents with different specialties (say, an HR assistant, a coder agent, and a data analyst agent) and, using AGNTCY standards, have them discover each other, share tasks, and coordinate workflows even if they were built by different vendors. Some early demonstrations (and code releases in 2025) show agents registering in a common directory and then composing their abilities to solve multi-step problems, illustrating the potential of this interoperability. The design goals of AGNTCY heavily emphasize extensibility and community governance. Rather than dictating a single closed protocol, the initiative encourages developers to extend the specs and contribute new ideas, hoping to avoid fragmentation by uniting efforts under one open umbrella. Its backers liken its importance to that of fundamental internet protocols, arguing that a similarly neutral and extensible standard for AI agents will unlock innovation across industries.
As of early 2025, AGNTCY is still ramping up; its initial specifications are available and open-source, but it acknowledges that broad adoption is crucial for success. In terms of limitations, AGNTCY’s challenges echo those of any nascent standard: it currently lacks widespread implementation, and there is a risk of competing protocols (like those above) dividing the community. The effort needs to onboard many AI platforms and toolmakers to truly become the “standard” – in other words, it faces an uphill battle for critical mass, where if not enough parties adopt it, the goal of universal agent interoperability could fail due to multiple incompatible ecosystems. Additionally, while AGNTCY covers discovery and connection, it is still developing the full “coordination logic” for complex workflows (their roadmap includes stages for orchestration and monitoring of multi-agent systems). This means that questions of how agents plan joint tasks, agree on protocols for negotiation, or maintain long-term collaborations may require further conventions on top of AGNTCY’s base layer.
@@ 3.2   Why Coral Protocol?
The AI agent ecosystem is rapidly evolving, with many projects striving to enable interoperable multi-agent systems. Google’s Agent-to-Agent (A2A) protocol and the open-source Agent Network Protocol (ANP) introduce common schemas and decentralized identity layers to let agents “speak” a shared language across frameworks. Meanwhile, initiatives like Cisco’s AGNTCY and MIT’s NANDA articulate grand visions of a decentralized Internet of Agents. Even commercial platforms (e.g. Synergetics.ai) are building agent orchestration features. These efforts demonstrate the community’s recognition that heterogeneous agents must collaborate. However, no single approach so far covers all the requirements for an open, dynamic agentic economy. A2A, for example, provides framework-agnostic messaging and task negotiation, but focuses on point-to-point calls rather than large-scale team formation. ANP offers strong cryptographic identity and peer-to-peer protocol negotiation, yet leaves out built-in incentives or easy team assembly. NANDA and AGNTCY set out an ambitious blueprint (registries, discovery, governance), but remain largely conceptual with no turnkey payment or task-allocation protocols. In short, existing initiatives tackle pieces of the problem, but gaps remain in secure team-building, economic incentives, and a unified interoperability layer.
Coral Protocol fills these gaps by synthesizing best practices from prior work into a complete, open-stack solution. Its design is vendor-neutral and modular, so that any agent framework can plug in without rewriting core logic. Like A2A, Coral defines a “common language” of message types and metadata that any agent can use, but it does so in a modular layer architecture that vendors can extend. Coral also incorporates a full payment and incentive layer: every agent and tool can register offers in a shared market and receive on-chain micro-payments for services. This economic fabric is inspired by NANDA’s insight that “well-designed incentive mechanisms…enable entirely new forms of collaboration…unlocking substantial economic opportunities.”. Crucially, Coral natively supports secure team formation: agents use decentralized identifiers (DIDs) and multi-party signatures (inspired by ANP’s use of W3C DIDs) to form consortia that can jointly bid on tasks or verify joint outputs. In practice, this means Coral can orchestrate many agents into ad-hoc coalitions for large jobs, with all communications end-to-end encrypted and group actions cryptographically authorized (similar to ANP’s meta-protocol for self-organizing collaboration networks).
Specifically, Coral Protocol offers:
  • Vendor-neutral interoperability: Coral reuses ideas from A2A’s agent cards and ANP’s schemas to ensure any compliant agent can join the network. Agents advertise their capabilities and protocols in a common format, so ecosystems are not siloed. As the A2A spec notes, agents gain a “common language – irrespective of the framework or vendor”, and Coral builds on this by publishing open specifications (JSON schemas, DID methods, etc.) under a neutral governance model.
  • Secure team creation: Coral adds a crypto-based team layer. Groups of agents can form trusted task forces. A team can collectively sign an agreement or share secret state, ensuring that multi-agent solutions are tamper-resistant and that agents cannot defect from a collaborative task without consensus. This solves a limitation of earlier proposals, which largely considered only single-agent endpoints.
  • Integrated payments and incentives: Unlike most protocols, Coral has a built-in token economy. Tasks, data, and services are all commoditized: agents stake tokens to bid on tasks or vouch for each other, and completed tasks automatically trigger crypto-payments. This aligns economic incentives at protocol level. As noted in analyses of NANDA, adding payment mechanisms can shift agent collaboration “beyond incremental efficiency” and unlock new marketplaces. Coral’s payment layer also funds network operations (for example, decentralized oracles and registries) in a self-sustaining way.
  • Dynamic multi-agent workflows: Coral is optimized for complex, many-to-many coordination, not just one-to-one requests. Agents can advertise skills, delegate subtasks, and negotiate task splits on the fly. Drawing on the A2A notion of agents exchanging capabilities and negotiating interaction patterns and on ANP’s meta-protocol for “self-negotiating collaboration networks”, Coral’s runtime supports emergent agent workflows. For example, when a high-level goal arrives, Coral’s discovery and directory services can assemble the optimal team of experts, multicast the goal, collect results, and merge them—all automatically.
  • Open, decentralized “Internet of Agents” vision: At its core, Coral seeks to realize an open agentic web. It embraces public interoperability standards so that any party can build compatible agents. It is vendor-neutral in governance, avoiding any single corporate control. This matches the articulated Internet-of-Agents ethos – “similar to how the original internet wasn’t built by a single entity” – with collaboration spanning organizations. By integrating features from protocols like A2A and ANP into an open protocol stack, while also addressing missing pieces (payments, team logic), Coral provides the missing substrate for a truly decentralized agent network.
In summary, Coral Protocol is not just another point solution; it is intended as the unifying framework for the Internet of Agents. By bringing together modular messaging, secure group orchestration, incentive mechanisms, and open standards, Coral overcomes the siloed approaches of past efforts. It thereby paves the way for a vibrant ecosystem where heterogeneous AI agents and humans can compose, compete, and cooperate globally. In this way, Coral realizes a vendor-neutral, scalable, and economic agent fabric that previous efforts only sketched out, moving us closer to an open, decentralized Internet of Agents.

### v2

# Coral Protocol Open Infrastructure Connecting The Internet of Agents

@@ Abstract
Coral Protocol is an open and decentralized collaboration infrastructure that enables communication, coordination, trust and payments for The Internet of Agents. It addresses the growing need for interoperability in a world where organizations are deploying multiple specialized AI agents that must work together across domains and vendors. As a foundational platform for multi-agent AI ecosystems, Coral establishes a common language and coordination framework allowing any agent to participate in complex coordinations with others. Its design emphasizes broad compatibility, security, and vendor neutrality, ensuring that agent interactions are efficient and trustworthy. In particular, Coral introduces standardized messaging formats for agent communication, a modular coordination mechanism for orchestrating multi-agent tasks, and secure team formation capabilities for dynamically assembling trusted groups of agents. Together, these innovations position Coral Protocol as a cornerstone of the emerging “Internet of Agents,” unlocking new levels of automation, collective intelligence, and business value through open agent collaboration.
@@ 1   Introduction
  • Dynamic Agent Discovery and Capability Registration: Agents can advertise their capabilities and discover others through standardized mechanisms, allowing seamless composition of multi-agent coordinations without hardcoded integrations or platform-specific logic.
  • Secure Team Formation and Task Execution Coordination: Coral enables the on-demand assembly of agent teams with authenticated identities, assigned roles, and controlled data access. Teams can collaboratively execute complex tasks while preserving privacy, trust, and auditability via a secure coordination layer and optional blockchain-based logging.
@@ 2   Enabling Concepts
In this section, we review several key concepts that underpin Coral Protocol’s vision of interoperable AI agents. We cover recent progress in large language models, autonomous AI agents and their use of tools, the emergence of model context protocols, and the advent of multi-agent AI systems. Together, these advances set the stage for why a unifying interoperability protocol is needed.
@@ 2.4   Model Context Protocol (MCP)
2.5   Multi AI Agent Systems
@@
There are several motivations for using multiple AI agents together. One is specialization: one agent might be an expert in math, another in coding, another in interacting with humans. By having them communicate, each sub-problem can be handled by the best-suited agent. For example, in a software development assistant, one agent could generate code while another agent reviews it for errors – akin to a pair programming scenario. Another motivation is parallelism: agents can work on different subtasks simultaneously and then share results, speeding up complex tasks. Yet another is emergent behavior: sometimes groups of agents can exhibit intelligent behaviors that single agents cannot, by bouncing ideas off each other or negotiating. A recent study by Tran et al. (2025) notes that LLM-based multi-agent systems enable "groups of intelligent agents to coordinate and solve complex tasks collectively at scale, transitioning from isolated models to collaboration-centric approaches" [19]. In other words, we’re beginning to move from thinking about one AI in isolation to networks of AIs collaborating, which some see as steps toward an "artificial collective intelligence."
@@ 2.6   Multi AI Agent Collaboration
Several proof-of-concept systems have explored multi-agent collaboration. The CAMEL framework (2023) showed that two LLM agents can role-play as a "user" and an "assistant" to iteratively solve a task together, essentially allowing the AI to have a conversation with itself from different perspectives to refine a solution. Microsoft’s AutoGen [21] provides a programming framework for composing *conversational* interactions among multiple agents (and humans), so that one can build execution plans where agents ask each other for help [21]. For instance, one agent could be tasked with decomposing a problem into steps, then delegating each step to other specialized agents, and finally aggregating the results. Experiments with AutoGen found that such setups can handle complex queries in domains like coding or decision-making more effectively than a single agent working alone [21]. Another striking example is the Generative Agents simulation mentioned earlier, where 25 agents inhabited a small virtual town – here the focus was on agents having social interactions with each other (e.g. sharing information, forming plans to organize a party) and it demonstrated that coherent multi-agent dynamics can emerge from relatively simple principles plus consistent language-model reasoning [16].
@@
This is where an interoperability framework becomes crucial: it can provide a structured medium for agent interaction (e.g., a shared memory or a common set of message types). The Coral Protocol is positioned in this landscape as an infrastructure to help organize and facilitate these multi-agent interactions. It extends ideas from MCP so that agents not only access data/tools uniformly, but also discover and communicate with each other in a standardized way. By having a common protocol, an "agent society" can form where each agent knows how to announce its capabilities, listen for requests, and share results in a mutual language. This would allow, for example, an agent built by Company A to collaborate with an agent from Company B if both speak Coral/MCP, much as devices on the internet interoperate via TCP/IP. In summary, multi-agent AI systems are an exciting and fast-evolving area — moving from single, siloed AI assistants towards networks of cooperating agents. The recent progress in this domain (just in the past couple of years) underscores the need for interoperability solutions: to manage complexity, avoid reinventing integration logic, and unlock the full potential of collective AI intelligence. Coral Protocol’s goal of facilitating agent interoperability directly addresses this need, aiming to enable robust, framework-agnostic collaboration among AI agents in the coming "society of AI agents."
@@ 2.8   Blockchains and Secure Payments
Blockchains provide a decentralized, tamper-evident ledger for recording all financial exchanges among agents. By design, every transaction written to the chain is immutable and publicly verifiable, ensuring reliable auditability. In Coral Protocol, agent payments (and any inter-agent transfers) are committed to the blockchain, creating an indelible audit trail of who paid whom and when. This means that all payments for services are transparently recorded and cannot be forged or altered after the fact.
@@ 3.1.1 Agent-to-Agent (A2A) Protocol
Google introduced the Agent-to-Agent (A2A) protocol, an open, vendor-neutral standard designed to enable seamless communication and collaboration among AI agents across various platforms and ecosystems [18]. A2A focuses on task-oriented interactions, providing standardized task objects with clearly defined lifecycles, and supporting the exchange of task-related artifacts to facilitate collaborative task executions. Built on established web standards like HTTP, JSON-RPC, and Server-Sent Events (SSE), the protocol ensures compatibility with existing technologies while offering enterprise-grade security through robust authentication and authorization aligned with OpenAPI. Additionally, A2A supports dynamic discovery of agent capabilities via JSON-based "agent cards" and handles both short-term tasks and long-running processes with real-time feedback. Google’s collaboration with over 50 industry partners—including Salesforce, SAP, and Workday—highlights the broad industry commitment toward achieving standardized agent interoperability.
@@
In practice, A2A is poised to facilitate complex task executions in both enterprise and open-source contexts. In enterprises, multiple AI agents could coordinate processes across business platforms that typically don’t talk to each other today. For example, an agent integrated with a CRM like Salesforce might invoke another agent on an ERP system such as SAP to automatically fulfill an order or update financial records [6]. Organizations are exploring scenarios like a sales assistant agent requesting a finance agent to generate a pricing quote, or a customer support chatbot querying an inventory management agent for stock availability – all through A2A’s standardized agent-to-agent calls without human middleware [8]. Because A2A is an open and framework-neutral protocol, it also lends itself to open-source innovation. Developers and researchers can orchestrate modular AI agents (for instance, linking a data analysis agent with a visualization agent) using A2A as the lingua franca for inter-agent dialogue. Indeed, popular toolkits like LangChain have begun integrating A2A, allowing independent or specialized agents to be composed into larger AI workflows in research and development settings.
@@
Despite its promise, the A2A protocol has some limitations that motivate the development of more comprehensive multi-agent frameworks. Notably, A2A focuses on the mechanics of messaging and doesn’t define a shared ontology or common knowledge base for semantics – agents still must understand the content of messages (often plain text or JSON) based on their own models or agreements, which can limit out-of-the-box interoperability in complex domains. The protocol is also in an early stage (initially released as a draft specification in 2025), so its extensibility to all possible use cases is still evolving and industry adoption is just beginning. Moreover, A2A by itself addresses communication rather than higher-level cognition or coordination among agents. It enables agents to talk but does not inherently provide collective planning, sophisticated negotiation, or “hive-mind” reasoning capabilities. Implementing such intelligence requires additional layers on top of A2A – for example, orchestration frameworks like Google’s Agent Development Kit (ADK) are intended to manage task executions and decision logic among multiple agents using the A2A channel. In summary, A2A lays an important foundation for agent interoperability, but it remains a low-level protocol; achieving robust multi-agent collaboration will also require tackling semantic standards, advanced coordination strategies, and other gaps that A2A alone does not fill.
@@ 3.1.3 AGNTCY (Internet of Agents Initiative)
Together, these pieces enable what AGNTCY envisions: for example, an organization could deploy multiple AI agents with different specialties (say, an HR assistant, a coder agent, and a data analyst agent) and, using AGNTCY standards, have them discover each other, share tasks, and coordinate task executions even if they were built by different vendors. Some early demonstrations (and code releases in 2025) show agents registering in a common directory and then composing their abilities to solve multi-step problems, illustrating the potential of this interoperability. The design goals of AGNTCY heavily emphasize extensibility and community governance. Rather than dictating a single closed protocol, the initiative encourages developers to extend the specs and contribute new ideas, hoping to avoid fragmentation by uniting efforts under one open umbrella. Its backers liken its importance to that of fundamental internet protocols, arguing that a similarly neutral and extensible standard for AI agents will unlock innovation across industries.
As of early 2025, AGNTCY is still ramping up; its initial specifications are available and open-source, but it acknowledges that broad adoption is crucial for success. In terms of limitations, AGNTCY’s challenges echo those of any nascent standard: it currently lacks widespread implementation, and there is a risk of competing protocols (like those above) dividing the community. The effort needs to onboard many AI platforms and toolmakers to truly become the “standard” – in other words, it faces an uphill battle for critical mass, where if not enough parties adopt it, the goal of universal agent interoperability could fail due to multiple incompatible ecosystems. Additionally, while AGNTCY covers discovery and connection, it is still developing the full “coordination logic” for complex task exeuctions (their roadmap includes stages for orchestration and monitoring of multi-agent systems). This means that questions of how agents plan joint tasks, agree on protocols for negotiation, or maintain long-term collaborations may require further conventions on top of AGNTCY’s base layer.
@@ 3.1.4 NANDA (Networked Agents and Decentralized AI)
3.1.6 IBM watsonx Orchestrator
IBM introduced watsonx Orchestrator [^3] Also marketed as IBM watsonx Orchestrate. as an enterprise-focused platform for multi-agent workflow automation and coordination [7]. Rather than defining just a communication protocol, watsonx Orchestrator delivers an end-to-end environment to build, run and manage AI agents that can collaborate across diverse business applications and data sources. Key capabilities announced in 2025 include a no-code Agent Builder (enabling custom AI agents to be created in under five minutes) and an Agent Catalog listing over 150 pre-built agents and tools from IBM and partners [7]. IBM provides domain-specialized agents (e.g. for HR, procurement, sales) with ready-made skills, alongside utility agents for tasks like web research and calculations, to jump-start common enterprise use cases [7]. These agents—and any custom or third-party agents—are orchestrated through a unified interface, with a central AI-driven planner intelligently routing tasks to the appropriate agent, tool, or human stakeholder as needed. Powered by large language models (e.g. IBM’s Granite series) for reasoning and planning, the Orchestrator acts as a “multi-agent supervisor, router and planner,” autonomously coordinating agents to complete complex workflows while abstracting away the complexity for end users [7].
A core strength of watsonx Orchestrator is its deep integration into enterprise IT ecosystems. The platform comes pre-integrated with 80+ enterprise applications and services (Salesforce, SAP, Workday, Microsoft tools, etc.), allowing AI agents to directly interact with existing software stacks for actions like database updates, form submissions, or ticket handling [7]. This out-of-the-box connectivity, combined with support for hybrid cloud and on-premise deployments, reflects IBM’s emphasis on immediate business value—agents can “plug into” legacy systems and cloud services without extensive custom integration. In contrast to pure protocol efforts, IBM’s solution provides built-in workflow orchestration: it not only enables messaging between agents but can automatically sequence multi-step processes (planning, invoking one agent after another, aggregating results, handling exceptions) to fulfill a high-level task. Enterprise-grade features such as agent observability and governance are also included [7]. For example, watsonx Orchestrator supplies monitoring dashboards and guardrails to track agent performance, data usage, and compliance, leveraging IBM’s AI governance frameworks for risk management. This focus on management and oversight addresses concerns like reliability, security and accountability which are critical in corporate settings.
Although IBM’s offering is a proprietary platform, it is designed with interoperability in mind. The Orchestrator’s extensible architecture allows agents built on various frameworks (including open-source agent libraries like LangChain, LangGraph, etc.) to be onboarded via standardized interfaces. IBM has introduced an Agent Connect developer framework that exposes familiar REST/JSON and chat-based APIs for external agents to communicate within the Orchestrator environment. Notably, watsonx Orchestrator also embraces emerging agent communication standards: it supports Anthropic’s Model Context Protocol (MCP) for tool and data access interoperability, and IBM has open-sourced an Agent Collaboration Protocol (ACP) to facilitate agent-to-agent messaging within its ecosystem [7]. By aligning with standards like MCP (and potentially bridging to protocols like A2A), IBM aims to avoid isolating its agents from the wider “internet of agents.” In essence, the platform implements a superset of agent communication functionality — a practical application layer that sits atop low-level protocols, abstracting them into a cohesive workflow engine.
@@ 3.2   Why Coral Protocol?
The AI-agent ecosystem is rapidly diversifying. Google’s Agent-to-Agent (A2A) protocol and the open-source Agent Network Protocol (ANP) introduce common schemas and decentralized identity layers so agents can “speak” a shared language. IBM’s watsonx Orchestrator demonstrates an enterprise-grade, no-code environment that plans and supervises multi-agent workflows while offering deep, out-of-the-box integrations with systems like Salesforce, SAP, and Workday. Meanwhile, consortium efforts such as Cisco’s AGNTCY and MIT’s NANDA outline ambitious blueprints for a decentralized Internet of Agents, and commercial stacks like Synergetics.ai add wallet-centric marketplaces.
Yet each of these initiatives addresses only a slice of the full problem space.
  • A2A excels at framework-agnostic message exchange but is point-to-point and leaves incentives and team logic to higher layers.
  • ANP provides strong DID-based identity and peer discovery, yet omits built-in payment rails or economic coordination.
  • IBM watsonx Orchestrator offers AI-driven planning and governance dashboards, but—as a proprietary platform—cannot guarantee vendor-neutral participation or tokenised incentives beyond IBM’s ecosystem.
  • AGNTCY and NANDA articulate open registries and governance but remain largely conceptual, with limited turnkey tooling for secure team assembly or on-chain value transfer.
Coral Protocol closes these gaps by combining the strongest ideas from prior work into a cohesive, open stack:
  • Vendor-neutral interoperability and extensibility. Coral re-uses A2A-style capability cards and ANP-inspired DID schemas, allowing any agent framework (including watsonx agents) to plug in without rewriting core logic.
  • Secure team formation. Agents cryptographically bind into ad-hoc coalitions using multi-party signatures and shared DIDs, enabling large-scale task forces that no point solution currently supports.
  • Integrated payments and incentives. Every agent or tool can publish services in a shared market and receive on-chain micro-payments, turning collaboration into an economy—something absent from A2A, ANP, and watsonx Orchestrator alike.
  • Dynamic multi-agent teams. Discovery, delegation, and result aggregation are baked into the runtime so that complex objectives can be decomposed across many agents—extending watsonx-style orchestration beyond a single vendor’s boundary.
  • Open, decentralized governance. Specifications, reference code, and registries are stewarded in a neutral foundation, ensuring that no single corporation controls the evolving “Internet of Agents” fabric.
In short, Coral Protocol is not merely another niche solution—it is designed as the unifying substrate that weaves together messaging, security, economic alignment, and open standards, enabling heterogeneous AI agents and humans to cooperate at global scale.